# Search-Based Credit Assignment for Offline Preference-Based Reinforcement Learning

### v2

@@ Introduction
$\mathcal{L}_{\text{CE}}=-\mathbb{E}_{(\tau^{0},\tau^{1},l)\sim\mathcal{D}}\Big{[}$ $(1-l)\log P_{\psi}[\tau^{0}\succ\tau^{1}]$
$+\;l\log P_{\psi}[\tau^{1}\succ\tau^{0}]\Big{]}.$ (3)
@@
To answer research question (1), we compare SPW with five offline PbRL baselines. Markovian Reward (MR), Preference Transformer (PT) (Kim, Han, and Park 2023), Offline PbRL (OPRL) (Shin, Dragan, and Brown 2023), Inverse Preference Learning (IPL) (Hejna and Sadigh 2024), and Listwise Reward Estimation (LiRE) (Choi, Park, and Lee 2024). The experimental results are shown in Table 1. MR method estimates reward with a simple Markovian MLP and performs poorly in most environments, reflecting its limited expressiveness. PT replaces MLP with a transformer to capture temporal dependencies, improving tasks such as sweep, handle-pull-side, but it underperforms on others, probably due to increased model complexity. OPRL leverages model ensembles for active query selection, but yields only marginal improvements over MR. IPL bypasses reward modeling and directly learns policies from preferences, but consistently lags behind, underscoring the challenges of this strategy. LiRE reformulates preferences as listwise rankings and replaces the BT model with a linear model, but still falls short of SPW.
@@
This section investigates the influence of the softmax temperature $\tau$ on policy success rate. Figure 5 reports performance when $\tau\!\in\!\{0.5,0.7,1,2,\infty\}$ while the total number of preference feedbacks is fixed at 200. Larger $\tau$ values produce smoother weight distributions. When $\tau\!\rightarrow\!\infty$, the weights become uniform and our method degenerates to MR. The best performance is obtained at $\tau=0.7$. As $\tau$ increases, the impact of the importance weights diminishes. Credit assignment within trajectories deteriorates, and the results degrade. When $\tau=\infty$, no weighting is applied. The outcome is clearly worse than the weighted variants. This further confirms that our approach improves reward credit assignment through preference‑trajectory weighting.
@@
After computing the weights between preference–trajectory transitions and expert trajectories, we implement the redistribution strategy from Hindsight PRIOR (Verma and Metcalf 2024). Concretely, we first compute the per‑step predicted rewards $\hat{r}_{\psi}(s_{t},a_{t})$ for each trajectory segment, then sum them to obtain the segment‑level reward: $R_{\psi}(\tau)=\sum_{t\in\tau}\hat{r}_{\psi}(s_{t},a_{t}).$ This total reward is redistributed to each step according to its learned weight $w_{t}$. The network is trained with an additional mean‑squared‑error (MSE) loss: $\mathcal{L}{\text{prior}}=\mathrm{MSE}\left(\hat{r}_{\psi}(s_{t},a_{t}),,w_{t},R_{\psi}(\tau)\right)$. Each step’s target reward becomes its weight $w_{t}$ times the segment return. We denote this weighted‑redistribution variant as RD.

### v3

@@ Introduction
$\mathcal{L}_{\text{CE}}=-\mathbb{E}_{(\sigma^{0},\sigma^{1},l)\sim\mathcal{D}_{pref}}\Big[$ $(1-l)\log P_{\psi}[\sigma^{0}\succ\sigma^{1}]$
$+\;l\log P_{\psi}[\sigma^{1}\succ\sigma^{0}]\Big].$ (3)
@@
To answer research question (1), we compare SPW with five offline PbRL baselines. Markovian Reward (MR), Preference Transformer (PT) (Kim, Han, and Park 2023), Offline PbRL (OPRL) (Shin, Dragan, and Brown 2023), Inverse Preference Learning (IPL) (Hejna and Sadigh 2024), and Listwise Reward Estimation (LiRE) (Choi, Park, and Lee 2024). The experimental results are shown in Table 1. MR method estimates reward with a simple Markovian MLP and performs poorly in most environments, reflecting its limited expressiveness. OPRL leverages model ensembles for active query selection, but yields only marginal improvements over MR. PT replaces MLP with a transformer to capture temporal dependencies, improving tasks such as peg-unplug-side, sweep-into, but it underperforms on others, probably due to increased model complexity. IPL bypasses reward modeling and directly learns policies from preferences, but consistently lags behind, underscoring the challenges of this strategy. LiRE reformulates preferences as listwise rankings and replaces the BT model with a linear model, but still falls short of SPW.
@@
This section investigates the influence of the softmax temperature $\tau$ on policy success rate. Figure 5 reports performance when $\tau\in\{0.5,0.7,1,2,\infty\}$, while the total number of preference feedbacks is fixed at 200.
Larger $\tau$ values produce smoother weight distributions. When $\tau\to\infty$, the weights become uniform and our method degenerates to MR. The best performance is obtained at $\tau=0.7$. As $\tau$ increases, the impact of the importance weights diminishes, leading to poorer credit assignment within trajectories and degraded results. When $\tau=\infty$, no weighting is applied, and the outcome is clearly worse than the weighted variants.
This further confirms that our approach improves reward credit assignment through preference-trajectory weighting.
@@
After computing the weights between preference–trajectory transitions and expert trajectories, we implement the redistribution strategy from Hindsight PRIOR (Verma and Metcalf 2024). Concretely, we first compute the per-step predicted rewards $\hat{r}_{\psi}(s_{t},a_{t})$ for each trajectory segment, then sum them to obtain the segment-level reward: $R_{\psi}(\sigma)=\sum_{t\in\sigma}\hat{r}_{\psi}(s_{t},a_{t}).$ This total reward is redistributed to each step according to its learned weight $w_{t}$. The network is trained with an additional mean-squared-error (MSE) loss. Each step’s target reward becomes its weight $w_{t}$ times the segment return. We denote this weighted-redistribution variant as RD. The prior loss is defined as:
$\mathcal{L}_{\mathrm{prior}}=\mathbb{E}_{\sigma\sim\mathcal{D}_{pref}}\left[\sum_{t\in\sigma}\big(\hat{r}_{\psi}(s_{t},a_{t})-w_{t}\,R_{\psi}(\sigma)\big)^{2}\right].$